Fermi National Laboratory
 

Accelerator Update for July 19, 2006 - July 21, 2006

 

Wednesday July 19

The day shift began with Operations monitoring store 4839 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, with the NuMI horn cooling work continuing, and with MiniBooNE taking beam.

A Pbar expert reported at 8:57 AM that he had valved in both cooling pumps at Pbar and the system temperatures were coming down.

At 9:05 AM, Operations received a report of a LCW leak in the lower Linac in the area of the Neutron Therapy Facility (NTF) water skid. A Linac expert investigated and reported that the leak was on the Linac test station. He will investigate further.

Operations established beam to the Meson MTest experiment at 9:16 AM.

A Pbar expert began a target scan at 9:19 AM.

Booster experts began tuning for higher intensities at 10:09 AM.

Operations began preparing the Recycler at 11:19 AM to stash antiprotons.

At 12:17 PM, Operations reported that the Recycler low-level RF went into a strange state. It wasn’t responding to commands. Operations halted the stash and resumed stacking.

MTest resumed taking beam 12:22 PM.

At 12:55 PM, Booster experts began working on a RF station (BRF17). This held off beam for fifteen minutes.

Operations reported at 1:36 PM that the heat exchanger for the Booster LCW system was clogged. Technicians brought the backup heat exchanger online and then began cleaning the main exchanger.

The Recycler began preparing to stash at 3:07 PM.

At 3:47 PM, Operations reported that store 4839 aborted due to a Tevatron (TeV) RF station (TRFF2) trip. Operators and experts investigated.

The stash ended and MTest resumed taking beam at 3:54 PM.

Operations prepared D0 for a collision hall access at 4:10 PM.

TeV RF experts reported at 5:26 PM that they were working on problems with TRF2, 3, and 4.

Operations prepared the CDF collision hall for access at 5:53 PM.

TeV RF experts reported that they were done with their work at 6:05 PM.

D0 came out of access at 6:41 PM.

CDF came out of access at 7:01 PM.

Operations put the TeV through a dry squeeze at 7:43 PM.

The MTest experiment finished using the beam at 8:27 PM.

Operations put the TeV into shot setup at 8:28 PM.

MI and Instrumentation experts began working on the MI BPMs at 8:33 PM.

At 8:34 PM, the Recycler damper tripped off and caused a transverse beam emittance blow up. An expert came in and worked at getting the damper back on to resume normal beam cooling.

MI experts began a BPM study at 9:02 PM.

Operations established store 4841 at 11:44 PM with an initial luminosity of 125.1E30.

The Recycler began preparing to stash at 11:45 PM.

Thursday July 20

The midnight shift began with Operations monitoring store 4841, with the Recycler preparing to stash, and with MiniBooNE taking beam.

At 12:10 AM, Operations reported that the stash was on hold due to a sequencer problem. The problem was resolved and the stash ended.

At 4:18 AM, Operations reported that store 4841 aborted. A TeV sector F2 CAMAC crate power supply failure had caused the ramp to trip off.

Operations put the TeV through a dry squeeze at 5:53 AM.

At 6:53 AM, Operations reported a flash of light, a clap of thunder, and then the MI power supply loop and MI-40 kickers went down. Everything reset okay.

The Recycler began stashing at 6:58 AM. The stash ended and Pbar resumed stacking at 7:06 AM.

At 7:09 AM, the TeV pond pump (M:C4PNDP) at sector C4 tripped off. Operators could not restart the pump and called the Duty Mechanic. He managed to get the pump back on an hour later.

Operations reported at 8 AM that the Recycler had stashed antiprotons three times during the midnight shift.

Operations put the TeV into a wet squeeze around 8 AM.

At 8:52 AM, Operations reported that the temperature in the TeV sector E2 service building electronics room was 104 degrees F. The Duty Mechanic investigated. Operations put the shot setup on hold.

The Duty Mechanic installed a portable air conditioner at E2 at 9:15 AM.

Operations put the TeV through a dry squeeze at 9:57 AM.

Operations put the TeV back into shot setup at 10:27 AM.

Operations established store 4845 at 1:31 PM with an initial luminosity of 102.68E30.

MTest resumed taking beam at 1:54 PM.

At 2:09 PM, BRF7 tripped off with no status to indicate what was wrong. An operator contacted a RF expert.

The Recycler began preparing to stash at 3:13 PM. The stash ended and Pbar resumed stacking at 4:05 PM. MTest resumed taking beam.

The Recycler began preparing to stash at 8:24 PM. The stash ended and Pbar resumed stacking at 9:04 PM. MTest resumed taking beam.

Friday July 21

The midnight shift began with Operations monitoring store 4845 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

The Recycler began preparing to stash at 1:12 AM. The stash ended and Pbar resumed stacking about 1:45 AM.

The Recycler began preparing to stash at 5:42 AM. The stash ended and Pbar resumed stacking at 6:07 AM.

The Plans for Friday and the Weekend

The plans for today are to terminate the store at 2 PM and then allow CDF and D0 to access their collision halls. The plans for the rest of the day and the weekend are to stack, stash, and store.

Machine Reports

Linac

Last Week

  • Running on the H- Ion Source
  • Beam Current 34.5 ma
  • Beam positions have varied throughout the week
  • Downtime: 6.69 Hours with 42 Occurrences — 96.0% Availability

Major Sources of Downtime

  • 2.42 hr Q5BULK Supply Problem
  • 1,57 hr. Various LRF5 Problems
  • .93 hr. Various Haefely Problems

Booster

Last Week

  • Stacking: 10-11 Turns, ~3.8 to 4.2E12 per batch
  • MiniBooNE: 10-11 Turns
  • NuMI: Off

Studies

  • Tuning injection orbit, Transition, HOPS (tunes) — Efficiency
  • Preaccelerator Notching continues on $1D cycles
  • Worked on dampers and conducted quadrupole damper studies

Downtime: 6 hours, plus a Linac half hour access

  • 400MeV wire (MW3)
  • RF Miscellaneous

Next Week

  • Continue to work on Preacc notching $1D
  • Push 12 turns on stacking
  • Damper work – Quad damper

Main Injector

MI operation

  • Shot to the TeV
    • Pbar coalescing eff. 85 - 91%
  • Shot to the RR
    • Efficiency 90 - 100%
  • Stacking
    • Intensity @injection ~8.0E12
    • Intensity @ target ~7.2E12 (Acc. eff. ~90%)
    • Beam loading compensation at transition

BPM upgrade

  • Orbit closure study for reverse protons
  • MI60 North electronics installation — next week MI studies

BLM upgrade

  • MI30 electronics installation – complete

MI Studies

  • Aperture scans continue
  • Beta function measurements

Antiproton Source

Beam on target

  • 7.1E12 —> 7.5E12 => 5%

Beamline auto orbit correction (a.k.a. Overthruster)

  • 120 GeV BPM DAQ problems not resolved yet
  • Will attempt to run 120 GeV Overthruster with P1 BPMs masked
  • AP2 correction requires stable 120 GeV orbit
  • This will reduce the large variation in stacking rate
    • Best stacking rate will not change
    • Average stacking rate will increase

Debuncher cooling kicker repair

  • >25% of Horizontal cooling + 12.5% of Dp cooling
  • Will be repaired during DØ 2-shift access (within next two weeks)
  • This may be why the D-A transfer looks like a bottleneck

Fix stacktail cooling

  • Will allow the Accumulator to stack the increased flux when all of the above fixes are complete

Plans

  • Continue trouble-shooting 120 GeV beamline BPMs
    • Try to make Overthruster work without P1 BPMs
    • When fixed re-establish 120 GeV beamline and AP2 auto orbit correction
  • Stacktail adjustments to increase Ed
  • Repair Debuncher kicker tank 201(?)
  • Match ARF4 to MIRF energy and phase
  • Continue AP3 beamline adjustments

Recycler

Transverse emittance growth during momentum mining/extraction

  • Typical parabolic profile in time for exiting parcel emittances (flying wires)
  • Average emittance and the impact on luminosity (assuming identical coalescing efficiency and proton emittances, assuming identical H. and V. emittances):
    • 7% for store 4808 (average = 4.8 π-mm-mr)
    • 10% for store 4818 (average = 6.3 π-mm-mr)
  • In reality it matters less
    • Better coalescing
    • Lower growth in one plane

Current criteria for Tevatron shots

  • What do we ask the operators to do?
    • Less than 5 π-mm-mr transverse emittances (Schottky)
    • Between 55 and 60 eVs independent of intensity
  • Good middle-ground between efficiency and stability

Next

  • Recycler switch tree cross-talk with accumulator
  • Tune space exploration
  • Antiproton lifetime during aggressive cooling
  • RF tests on development system

Tevatron

Store Summary

Store

Avg Init. Lum (E30)

CDF Deliv’d

Lum

(nb-1)

Time (hr)

How did it end?

Comments

4844

91.3

3200+

19+

Ongoing

 

4841

125.1

1602

4.7

Abort

F2 CAMAC crate died

4839

102.3

3356

21.5

Abort

TRF3 trip in squeeze – TRF2 trip abort

4838

79.2

3396

33.7

Intentional

End of store studies

4835

56.6

2027

19.7

Quench

C17 Separator spark – TRF2 trip caused abort gap issues

4829

136

28.1

4 min

Abort

A2 LCW temperature trip

4828

91.7

3903

27.3

Intentional

Still in from previous week

Problems

  • RF trips
  • A17 Separator polarity during studies
  • Water Temperature problems
  • Vertical Flying Wire not working
  • Proton Loss during antiproton injection (minor)

Work Done this Week

  • Working point adjustment for losses and lifetimes
  • MI BPM work for Reverse injection tune up
  • Coupling studies
  • B48 separator voltage adjustment
  • D0 horizontal angle change
  • HEP tune Mult calibrations
  • Tune monitor studies
  • A0 abort position monitoring alarm enabled
  • IP run during stores

Work for Next Week

  • Work on antiproton loss at start of ramp
  • Close Pbar injection bump (reduce Proton loss)
  • Coupling measurements during shot setup with Eliana’s program
  • Routinely use Struck system to close Proton Injection
  • Check bumper timing
  • Test remote setting of Tel pulse length
  • Tune monitor
  • Testing Linux applications
  • Orbit stabilization on ramp

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/28/2006    email Fermilab